OpenStudy (anonymous):

Factoring out \(x-3\) from \[\frac{\frac{1}{x-4}+1}{x-3}\]. Basically, I'm working on some limit questions, and I need to factor out \(x-3\) from the base so I can determine if there is a VA or not. The limit is \(x->3\), which makes both the denominator and numerator zero. Therefore, I need to get \(x-3\) into the numerator somehow so I can cancel it with the \(x-3\) in the denominator.

OpenStudy (anonymous):

hey can you please type the question again it does not really make sense like that?

OpenStudy (anonymous):

oh i just got it my screen was not showing it well

OpenStudy (anonymous):

you have to first simplify the numerator

OpenStudy (anonymous):

so that becomes (x+5)/(x-4) and all of this is divided by x-3

OpenStudy (anonymous):

Wow, that was dumb! I even multiplied by \(x-4\) top and bottom before, and didn't see, somehow, that 1+ x-4 in the numerator became x-3 (messy writing ftw!).

OpenStudy (anonymous):

now you can write this as ((x+5)/(x-4)) * (x-3)

OpenStudy (anonymous):

this ends up being: \[\frac{ (x+5)(x-3) }{ (x-4) }\] You cant simplify it anymore!

OpenStudy (anonymous):

I mean there is nothing that can be cancelled out.

OpenStudy (anonymous):

I think it becomes: \[\frac{(x-3)}{(x-3)(x-4)}\]

OpenStudy (anonymous):

i think you are multiplying x-3 to the denominator which is a common mistake

OpenStudy (anonymous):

Because \[(\frac{1}{x-4}+1) * (x-4) = 1 + x-4\]

OpenStudy (anonymous):

For what it's worth, it gave me the right answer!